on average, junior girls own 10 pairs of shoes with a standard deviation of 1.5 shoes. the distribution of shoes follows an approximately normal curve. what is the probability of selecting a junior who owns 12 or more pairs of shoes?
Answer by stanbon(75887)

on average, junior girls own 10 pairs of shoes with a standard deviation of 1.5 shoes. the distribution of shoes follows an approximately normal curve. what is the probability of selecting a junior who owns 12 or more pairs of shoes?
------
z(12) = (12-10)/1.5 = 2/1.5 = 4/3
------
P(x >= 12) = P(z >= 4/3) = 0.0912
=========================
Cheers,
Stan H.
===============